OpenStudy (anonymous):

explain two ways to determine if an equation is linear. can someone please help!!!!!

OpenStudy (anonymous):

@DollyAcquah can you help me plese

OpenStudy (anonymous):

@Nurali can you help me please

OpenStudy (anonymous):

@KateGrace can you help me

OpenStudy (anonymous):

I'm sorry but linear algebra is not my strong suit!

OpenStudy (anonymous):

But I think one might be when the equation is y=mx+b

OpenStudy (anonymous):

thanks for trying to help

OpenStudy (anonymous):

@Destinymasha can you help please

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

@kategrace is right. A lineear equation means that when you put it on a grph it makes a straight line. Two type of equations are... slope-intercept form: y = mx +b (like she said) and... point-slope form: y - y1 = m(x - x1)

OpenStudy (kamibug):

hope that helps and answers your question! =)